# Through the River: Understanding the Benefit of Schedule-Free Methods for Language Model Training

### v1

@@ 1 Introduction
To support prolonged and flexible training, practitioners have adopted more adaptive scheduling strategies. One widely used approach is the warmup-stable-decay (WSD) schedule (Hu et al., 2024), which avoids committing to a fixed compute budget by maintaining a main “branch” with a constant learning rate (LR) and periodically branching into decaying LR trajectories to produce intermediate checkpoints—enabling flexible and continued training. Despite its advantages, WSD has notable limitations. A key challenge lies in evaluating the quality of the current model without explicitly entering the decay phase. This lack of visibility complicates decisions around checkpointing and training continuation, leading to uncertainty in training management (see Section 2.2).
One common workaround is to maintain a weight averaging, which improves generalization and provides more stable performance estimation. However, this comes at the cost of additional memory overhead and implementation complexity, especially in distributed training setups (see Section 2.3).
@@
  • We revisit the river-valley loss landscape and analyze two widely used strategies—WSD and weight averaging—through this lens, highlighting their respective strengths and limitations for scalable pretraining (Section 2).
  • We then focus on the SF method and show that it effectively follows the “river” structure of the loss landscape without requiring a decay phase or auxiliary averaging. This makes it particularly well-suited for continuously scaling training workloads (Section 3).
  • We analyze its training dynamics both theoretically and empirically. Our findings reveal that SF implicitly performs a form of weight averaging, without requiring additional memory. We also show that it operates at the Edge of Stability (Cohen et al., 2021) and derive its associated central flow (Cohen et al., 2025), providing a deeper understanding of its behavior (Section 4).
  • Based on these insights, we propose a refined version of the SF method that improves robustness to momentum parameters and scale better with large batch sizes—addressing key limitations of the original method (Section 5).
@@ 2.1 Backgrounds: Loss Landscape of Neural Networks
Figure 1: River-valley structure in a toy loss landscape. Contour plot of the objective defined in Section 4.1, illustrating the flat river direction and steep hill direction characteristic of the river-valley geometry.
@@
As its name suggests, a river-valley loss landscape resembles a winding ravine: steep “hill” walls flank a relatively flat “river” floor that snakes through parameter space. Wen et al. (2025) formalize this picture to interpret the warmup-stable-decay (WSD) LR schedule, arguing that large, noisy updates during the stable phase let SGD travel quickly downstream along the river, while the fast-decay phase pulls the iterate back to the bottom of valley. A concrete illustration is provided in Figure 1 (see also Section 4.1). Complementary evidence from Song et al. (2025) show that “effective” optimizer updates happen along the river: projecting out the high-curvature hill directions does not harm progress of learning, indicating that motion in the hill directions is often dispensable once the iterate is near the river.
@@
  • the river component, which measures progress along the low-curvature valley floor, and
@@ 2.4 Schedule-Free Methods
In this work, we focus on Schedule-Free AdamW (SF-AdamW), where $\Delta_{t}$ is computed using the RMSprop update along with a weight decay term. The full pseudocode is provided in Algorithm 1. Here, $\beta_{1}$ denotes the coefficient $\beta$ in (SF), and $\beta_{2}$ is the momentum parameter used in the second-moment of RMSprop.
@@ 3 Schedule-Free Optimizer as a Scalable Pretraining Method
In Section 2, we discussed the limitations of WSD and weight averaging as strategies for scalable pretraining. While WSD relies on a decay phase to achieve optimal performance, weight averaging avoids decay but incurs additional memory overhead. In this section, we empirically investigate the SF method as an alternative. We find it to be a strong candidate, as it requires neither decay nor extra memory.
@@
We use a 124M parameter LLaMA (Touvron et al., 2023a, b) style decoder-only transformer, trained with SF-AdamW using a warmup phase followed by a constant LR. The batch size is 0.5M tokens, and training is conducted on a 6B-token subset of SlimPajama (Soboleva et al., 2023), with the compute budget determined by the Chinchilla scaling rule (Hoffmann et al., 2022). We report validation loss in terms of perplexity. Additional results using a 124M parameter GPT-2 (Radford et al., 2019) style decoder-only transformer trained on the OpenWebText2 dataset (Gao et al., 2020) are provided in Appendix C. Full training details are available in Appendix B.
@@
As noted in Section 2, standard AdamW with a constant LR typically yields suboptimal performance, requiring a decay phase or weight averaging to reach better solutions. To test whether SF-AdamW exhibits similar behavior, we first perform a grid search to identify the best hyperparameters for both AdamW and SF-AdamW under constant LR (after warmup). In our setting, the optimal hyperparameters are $(\beta_{1},\beta_{2})=(0.9,0.95)$ with LR 1e-3 for AdamW, and $(\beta_{1},\beta_{2})=(0.95,0.99)$ with LR 2e-3 for SF-AdamW. Using these configurations, we train each model and periodically save checkpoints. At each checkpoint, we run a LR decay phase and evaluate the resulting loss. We also track the EWA of the ${\mathbf{x}}_{t}$ iterates throughout training. Results are shown in Figure 2.
@@
We next examine how closely the SF-AdamW trajectory follows the river in the loss landscape, building on the observation by Wen et al. (2025) that AdamW with a decaying LR converges toward the river. Specifically, we run a short decay phase of AdamW (linear decay from 1e-4 to 0) at each checkpoint from the SF-AdamW run described in the previous experiment. This decay phase—starting from a small LR—is designed to make minimal progress along the river direction while substantially reducing the hill component, pulling the iterate toward the valley floor. As a baseline, we apply the same procedure to AdamW. Results are shown in Figure 2.
We observe that applying a decay phase of AdamW to the SF-AdamW trajectory results in minimal additional loss reduction, in contrast to the AdamW trajectory where the decay phase leads to a sharp drop in loss. This suggests that SF-AdamW already closely tracks the river throughout training, eliminating the need for LR decay or weight averaging.
Figure 2: SF-AdamW closely follows the river, unlike AdamW. Left, Middle: While AdamW benefits from linear LR decay and EWA, SF-AdamW shows no improvement from either. Right: A short decay phase of AdamW (with linear LR decay from 1e-4 to 0) leads to a sharp loss drop for AdamW, but has no effect when applied to the SF-AdamW trajectory—suggesting that SF-AdamW already tracks the river throughout training (Figure 2).
@@
Figure 3: SF-AdamW with suboptimal momentum fails to follow the river. A short decay phase of AdamW applied to SF-AdamW checkpoints with $\beta_{1}\in\{0.1,0.5\}$ results in a sharp loss drop, unlike the case with $\beta_{1}=0.95$ (Section 3).
@@
To further investigate this gap, we repeat our experiment using SF-AdamW with suboptimal momentum $\beta_{1}\in\{0.1,0.5\}$. As before, we periodically save checkpoints and apply a short AdamW decay phase at each. Unlike the optimal case ($\beta_{1}=0.95$), we observe that the decay phase improves performance, suggesting that suboptimal momentum disrupts the optimizer’s ability to follow the river. Results are shown in Figure 3.
@@ 4.1 Warmup: A Toy Model
where ${\mathbf{w}}=(w^{(1)},w^{(2)})\in\mathbb{R}^{2}$. As shown in Figure 1, this objective exhibits a river-valley structure: the curve $w^{(1)}w^{(2)}=1$ forms a river, along which the loss slowly decreases as $w^{(1)}$ increases. We use this toy model to visualize the training dynamics of SF-AdamW in a controlled setting.
We run SF-AdamW with a constant LR, initializing at $\mathbf{x}_{1}=(2,2)$, fixing $\beta_{2}=0.99$ and varying $\beta_{1}\in\{0.1,0.5,0.9\}$. For each run, we track the iterates $({\mathbf{x}}_{t},{\mathbf{y}}_{t},{\mathbf{z}}_{t})$. Results are shown in Figure 4.
@@
This behavior aligns with Section 3, where we observed that SF-AdamW fails to follow the river under suboptimal momentum configurations in language model training. Notably, even when $\beta_{1}$ is suboptimal, the ${\mathbf{y}}_{t}$ iterates continue to track the river on average, despite exhibiting oscillations (we revisit the nature of this oscillation in Section 4.3). These observations suggest that the ${\mathbf{y}}_{t}$ sequence is more robust to $\beta_{1}$ and better aligned with the river geometry than the ${\mathbf{x}}_{t}$ iterates, making it a more reliable foundation for analyzing and guiding optimization in SF dynamics. We investigate this further in the context of language model training.
Figure 4: SF-AdamW on toy model. Left: The ${\mathbf{x}}_{t}$ iterates fail to follow the river for $\beta_{1}\in\{0.1,0.5\}$ (Section 3). Right: The ${\mathbf{y}}_{t}$ iterates oscillate around the river but track it reliably on average, even for suboptimal values of $\beta_{1}$ (Figure 5). As $\beta_{1}$ increases, the oscillations shrink.
@@ 4.2 Language Model Training
We evaluate the loss at the ${\mathbf{y}}_{t}$ iterates, as well as the EWA of ${\mathbf{y}}_{t}$, using the same experimental runs from Section 3 with $\beta_{1}\in\{0.1,0.5,0.95\}$. Results are shown in Figure 5. Notably, for suboptimal $\beta_{1}$, we observe that the loss at ${\mathbf{y}}_{t}$ is consistently lower than that at ${\mathbf{x}}_{t}$, showing that ${\mathbf{y}}_{t}$ more faithfully follows the river geometry and remains robust to suboptimal momentum settings. This mirrors our findings in the toy model analysis (Section 4.1), where ${\mathbf{x}}_{t}$ failed to follow the river under suboptimal momentum, while ${\mathbf{y}}_{t}$ continued to track it. Moreover, the EWA of ${\mathbf{y}}_{t}$ consistently achieves lower loss than the raw ${\mathbf{y}}_{t}$ iterates across all momentum configurations—unlike the ${\mathbf{x}}_{t}$ iterates, where EWA offers no benefit for $\beta_{1}=0.95$. It illustrates that the EWA of ${\mathbf{y}}_{t}$ consistently remains closer to the river compared to the vanilla ${\mathbf{y}}_{t}$ iterates. This observation parallels the oscillatory behavior of ${\mathbf{y}}_{t}$ in the toy model, where the EWA would more closely align with the underlying river-like geometry.
Figure 5: Performance of ${\mathbf{x}}_{t}$, ${\mathbf{y}}_{t}$, and the EWA of ${\mathbf{y}}_{t}$ under varying $\beta_{1}$. For suboptimal $\beta_{1}$, ${\mathbf{y}}_{t}$ outperforms ${\mathbf{x}}_{t}$, and across all momentum settings, the EWA of ${\mathbf{y}}_{t}$ achieves the lowest loss (Figure 5).
@@ 4.3 Schedule-Free Methods at the Edge of Stability
We return to the toy model from Section 4.1 to further analyze the optimization dynamics of the ${\mathbf{y}}_{t}$ iterates. Notably, the ${\mathbf{y}}_{t}$ sequence exhibits a period-two oscillation centered along the river (Figure 4). This behavior resembles the dynamics of full-batch GD at the Edge of Stability (EoS), where iterates oscillate along sharp directions while remaining globally stable. The EoS phenomenon was first identified empirically by Cohen et al. (2021) and has since been studied theoretically (Arora et al., 2022; Wang et al., 2022; Ahn et al., 2023; Damian et al., 2023; Song and Yun, 2023; Zhu et al., 2023). A key feature of this regime is that the sharpness—measured by the largest Hessian eigenvalue—stabilizes near the threshold $2/\gamma$.
@@
Notation.
Let $f({\mathbf{w}})$ denote the objective and $\mathbf{H}({\mathbf{w}})$ its Hessian at ${\mathbf{w}}$. For brevity, write $\mathbf{H}_{t}\coloneqq\mathbf{H}({\mathbf{w}}_{t})$. The largest eigenvalue $\lambda_{1}(\mathbf{H})$ is called the sharpness, and for a preconditioner $\mathbf{P}$, $\lambda_{1}(\mathbf{P}^{-1}\mathbf{H})$ is called the preconditioned sharpness.
We first the stability of Schedule-Free GD (SF-GD) on quadratic objectives, which serve as local Taylor approximations to neural network training losses. SF-GD is defined by (SF) with $\Delta_{t}\triangleq\nabla f({\mathbf{y}}_{t})$.
@@
Next, we extend the analysis to SF-PrecondGDW, defined by (SF) with $\Delta_{t}\triangleq\mathbf{P}^{-1}\nabla f({\mathbf{y}}_{t})+\lambda{\mathbf{y}}_{t}$, where $\mathbf{P}$ is a fixed, symmetric, and positive definite preconditioner, and $\lambda$ denotes the weight decay coefficient. This setting parallels the analysis of Adaptive Edge of Stability introduced by Cohen et al. (2022).
@@
Proofs of Propositions 4.1 and 4.2 are deferred to Section D.1.
SF-AdamW can be viewed as SF-PrecondGDW with with a slowly varying diagonal preconditioner $\mathbf{P}_{t}$. Hence, its stability is governed by the preconditioned sharpness $\lambda_{1}(\mathbf{P}_{t}^{-1}\mathbf{H}_{t})$. As shown in Figure 6, the preconditioned sharpness at ${\mathbf{y}}_{t}$ iterates equilibrates near the stability threshold in both the toy model and CIFAR-10 experiments—exhibiting a typical EoS behavior.
Figure 6: The ${\mathbf{y}}_{t}$ iterates of Schedule-Free methods operate at the Edge of Stability. Plots of (preconditioned) sharpness during full-batch training; dashed lines indicate stability thresholds. Left: Toy model trained using SF-AdamW with $(\beta_{1},\beta_{2})=(0.9,0.99)$. Middle, Right: Fully connected network trained on a 5k subset of CIFAR-10 using SF-GD with $\beta=0.9$ and SF-AdamW with $(\beta_{1},\beta_{2})=(0.95,0.99)$; (preconditioned) sharpness is evaluated at the $\mathbf{y}_{t}$ iterates (Figure 6).

### v4

@@ 1 Introduction
To support prolonged and flexible training, practitioners have adopted more adaptive scheduling strategies. One widely used approach is the warmup-stable-decay (WSD) schedule (Hu et al., 2024), which avoids committing to a fixed compute budget by maintaining a main “branch” with a constant learning rate (LR) and periodically branching into decaying LR trajectories to produce intermediate checkpoints—enabling flexible and continued training. Despite its advantages, WSD has notable limitations. A key challenge lies in evaluating the quality of the current model without explicitly entering the decay phase. This lack of visibility complicates decisions around checkpointing and training continuation, leading to uncertainty in training management (see Section˜2.2).
One common workaround is to maintain a weight averaging, which improves generalization and provides more stable performance estimation. However, this comes at the cost of additional memory overhead and implementation complexity, especially in distributed training setups (see Section˜2.3).
@@
  • We revisit the river-valley loss landscape and analyze two widely used strategies—WSD and weight averaging—through this lens, highlighting their respective strengths and limitations for scalable pretraining (Section˜2).
  • We then focus on the SF method and show that it effectively follows the “river” structure of the loss landscape without requiring a decay phase or auxiliary averaging. This makes it particularly well-suited for continuously scaling training workloads (Section˜3).
  • We analyze its training dynamics both theoretically and empirically. Our findings reveal that SF implicitly performs a form of weight averaging, without requiring additional memory. We also show that it operates at the Edge of Stability (Cohen et al., 2021) and derive its associated central flow (Cohen et al., 2025), providing a deeper understanding of its behavior (Section˜4).
  • Based on these insights, we propose a refined version of the SF method that improves robustness to momentum parameters and scale better with large batch sizes—addressing key limitations of the original method (Section˜5).
@@ 2.1 Backgrounds: Loss Landscape of Neural Networks
Figure 1: River-valley structure in a toy loss landscape. Contour plot of the objective defined in Section˜4.1, illustrating the flat river direction and steep hill direction characteristic of the river-valley geometry.
@@
As its name suggests, a river-valley loss landscape resembles a winding ravine: steep “hill” walls flank a relatively flat “river” floor that snakes through parameter space. Wen et al. (2025) formalize this picture to interpret the warmup-stable-decay (WSD) LR schedule, arguing that large, noisy updates during the stable phase let SGD travel quickly downstream along the river, while the fast‑decay phase pulls the iterate back to the bottom of valley. A concrete illustration is provided in Figure˜1 (see also Section˜4.1). Complementary evidence from Song et al. (2025) show that “effective” optimizer updates happen along the river: projecting out the high‑curvature hill directions does not harm progress of learning, indicating that motion in the hill directions is often dispensable once the iterate is near the river.
@@
  • the river component, which measures progress along the low‑curvature valley floor, and
@@ 2.4 Schedule-Free Methods
In this work, we focus on Schedule-Free AdamW (SF-AdamW), where $\Delta_{t}$ is computed using the RMSprop update along with a weight decay term. The full pseudocode is provided in Algorithm˜1. Here, $\beta_{1}$ denotes the coefficient $\beta$ in (SF), and $\beta_{2}$ is the momentum parameter used in the second-moment of RMSprop.
@@ 3 Schedule-Free Optimizer as a Scalable Pretraining Method
In Section˜2, we discussed the limitations of WSD and weight averaging as strategies for scalable pretraining. While WSD relies on a decay phase to achieve optimal performance, weight averaging avoids decay but incurs additional memory overhead. In this section, we empirically investigate the SF method as an alternative. We find it to be a strong candidate, as it requires neither decay nor extra memory.
@@
We use a 124M parameter LLaMA (Touvron et al., 2023a, b) style decoder-only transformer, trained with SF-AdamW using a warmup phase followed by a constant LR. The batch size is 0.5M tokens, and training is conducted on a 6B-token subset of SlimPajama (Soboleva et al., 2023), with the compute budget determined by the Chinchilla scaling rule (Hoffmann et al., 2022). We report validation loss in terms of perplexity. Additional results using a 124M parameter GPT-2 (Radford et al., 2019) style decoder-only transformer trained on the OpenWebText2 dataset (Gao et al., 2020) are provided in Appendix˜C. Full training details are available in Appendix˜B.
@@
As noted in Section˜2, standard AdamW with a constant LR typically yields suboptimal performance, requiring a decay phase or weight averaging to reach better solutions. To test whether SF-AdamW exhibits similar behavior, we first perform a grid search to identify the best hyperparameters for both AdamW and SF-AdamW under constant LR (after warmup). In our setting, the optimal hyperparameters are $(\beta_{1},\beta_{2})=(0.9,0.95)$ with LR 1e-3 for AdamW, and $(\beta_{1},\beta_{2})=(0.95,0.99)$ with LR 2e-3 for SF-AdamW. Using these configurations, we train each model and periodically save checkpoints. At each checkpoint, we run a LR decay phase and evaluate the resulting loss. We also track the EWA of the ${\mathbf{x}}_{t}$ iterates throughout training. Results are shown in Figure˜2.
@@
We next examine how closely the SF-AdamW trajectory follows the river in the loss landscape, building on the observation by Wen et al. (2025) that AdamW with a decaying LR converges toward the river. Specifically, we run a short decay phase of AdamW (linear decay from 1e-4 to 0) at each checkpoint from the SF-AdamW run described in the previous experiment. This decay phase—starting from a small LR—is designed to make minimal progress along the river direction while substantially reducing the hill component, pulling the iterate toward the valley floor. As a baseline, we apply the same procedure to AdamW. Results are shown in Figure˜2.
We observe that applying a decay phase of AdamW to the SF-AdamW trajectory results in minimal additional loss reduction, in contrast to the AdamW trajectory where the decay phase leads to a sharp drop in loss. This suggests that SF-AdamW already closely tracks the river throughout training, without requiring LR decay or weight averaging.
To further support this interpretation, we measure the loss along linear interpolations between the 2B and 2.5B token checkpoints under three training regimes: (1) AdamW with a constant learning rate (LR), (2) AdamW with a linear LR decay to zero, and (3) SF-AdamW with a constant LR. As shown in Figure 3, the resulting loss curves display qualitatively distinct behaviors: (1) shows a convex, valley-shaped profile, indicating oscillation across the valley; (2) shows a sharp, monotonic decline, consistent with a transition from the valley wall to the floor; and (3) shows a flat, slow decline, suggesting that the trajectory is already closely aligned with the valley floor. Notably, cases (1) and (2) replicate the loss profiles reported in Figure 7 of Wen et al. (2025), while (3) offers further evidence that SF-AdamW remains near the river throughout training.
Figure 2: SF-AdamW closely follows the river, unlike AdamW. Left, Middle: While AdamW benefits from linear LR decay and EWA, SF-AdamW shows no improvement from either. Right: A short decay phase of AdamW (with linear LR decay from 1e-4 to 0) leads to a sharp loss drop for AdamW, but has no effect when applied to the SF-AdamW trajectory—suggesting that SF-AdamW already tracks the river throughout training (Figure˜3).
Figure 3: Linear interpolation between training checkpoints. We evaluate the loss along linear interpolations $\alpha{\mathbf{w}}_{t_{1}}+(1{-}\alpha){\mathbf{w}}_{t_{2}}$, where $\alpha\in[0,1]$ and $t_{1}$, $t_{2}$ denote the 2B and 2.5B token checkpoints, respectively. We compare three training regimes: (1) AdamW with constant learning rate (LR), (2) AdamW with a linear LR decay to zero, and (3) SF-AdamW with constant LR. For all settings, the first 2B tokens are trained using either constant-LR AdamW (for 1 and 2) or constant-LR SF-AdamW (for 3). The resulting curves exhibit qualitatively distinct behaviors: convex (valley-shaped) for (1), sharp monotonic decay for (2), and flat, slowly declining loss for (3) (Figure˜3).
@@
Figure 4: SF-AdamW with suboptimal momentum fails to follow the river. A short decay phase of AdamW applied to SF-AdamW checkpoints with $\beta_{1}\in\{0.1,0.5\}$ results in a sharp loss drop, unlike the case with $\beta_{1}=0.95$ (Section˜3).
@@
To further investigate this gap, we repeat our experiment using SF-AdamW with suboptimal momentum $\beta_{1}\in\{0.1,0.5\}$. As before, we periodically save checkpoints and apply a short AdamW decay phase at each. Unlike the optimal case ($\beta_{1}=0.95$), we observe that the decay phase improves performance, suggesting that suboptimal momentum disrupts the optimizer’s ability to follow the river. Results are shown in Figure˜4.
@@ 4.1 Warmup: A Toy Model
where ${\mathbf{w}}=(w^{(1)},w^{(2)})\in\mathbb{R}^{2}$. As shown in Figure˜1, this objective exhibits a river-valley structure: the curve $w^{(1)}w^{(2)}=1$ forms a river, along which the loss slowly decreases as $w^{(1)}$ increases. We use this toy model to visualize the training dynamics of SF-AdamW in a controlled setting.
We run SF-AdamW with a constant LR, initialized at $\mathbf{x}_{1}=(2,2)$, fixing $\beta_{2}=0.99$ and varying $\beta_{1}\in\{0.1,0.5,0.9\}$. For each run, we track the iterates $({\mathbf{x}}_{t},{\mathbf{y}}_{t},{\mathbf{z}}_{t})$. Results are shown in Figure˜5.
@@
This behavior aligns with Section˜3, where we observed that SF-AdamW fails to follow the river under suboptimal momentum configurations in language model training. Notably, even when $\beta_{1}$ is suboptimal, the ${\mathbf{y}}_{t}$ iterates continue to track the river on average, despite exhibiting oscillations (we revisit the nature of this oscillation in Section˜4.3). These observations suggest that the ${\mathbf{y}}_{t}$ sequence is more robust to $\beta_{1}$ and better aligned with the river geometry than the ${\mathbf{x}}_{t}$ iterates, making it a more reliable foundation for analyzing and guiding optimization in SF dynamics. We investigate this further in the context of language model training.
Figure 5: SF-AdamW on toy model. Left: The ${\mathbf{x}}_{t}$ iterates fail to follow the river for $\beta_{1}\in\{0.1,0.5\}$ (Section˜3). Right: The ${\mathbf{y}}_{t}$ iterates oscillate around the river but track it reliably on average, even for suboptimal values of $\beta_{1}$ (Figure˜6). As $\beta_{1}$ increases, the oscillations shrink.
@@ 4.2 Language Model Training
We evaluate the loss at the ${\mathbf{y}}_{t}$ iterates, as well as the EWA of ${\mathbf{y}}_{t}$, using the same experimental runs from Section˜3 with $\beta_{1}\in\{0.1,0.5,0.95\}$. Results are shown in Figure˜6. Notably, for suboptimal $\beta_{1}$, we observe that the loss at ${\mathbf{y}}_{t}$ is consistently lower than that at ${\mathbf{x}}_{t}$, showing that ${\mathbf{y}}_{t}$ more faithfully follows the river geometry and remains robust to suboptimal momentum settings. This mirrors our findings in the toy model analysis (Section˜4.1), where ${\mathbf{x}}_{t}$ failed to follow the river under suboptimal momentum, while ${\mathbf{y}}_{t}$ continued to track it. Moreover, the EWA of ${\mathbf{y}}_{t}$ consistently achieves lower loss than the raw ${\mathbf{y}}_{t}$ iterates across all momentum configurations—unlike the ${\mathbf{x}}_{t}$ iterates, where EWA offers no benefit for $\beta_{1}=0.95$. It illustrates that the EWA of ${\mathbf{y}}_{t}$ consistently remains closer to the river compared to the vanilla ${\mathbf{y}}_{t}$ iterates. This observation parallels the oscillatory behavior of ${\mathbf{y}}_{t}$ in the toy model, where the EWA would more closely align with the underlying river-like geometry.
Figure 6: Performance of ${\mathbf{x}}_{t}$, ${\mathbf{y}}_{t}$, and the EWA of ${\mathbf{y}}_{t}$ under varying $\beta_{1}$. For suboptimal $\beta_{1}$, ${\mathbf{y}}_{t}$ outperforms ${\mathbf{x}}_{t}$, and across all momentum settings, the EWA of ${\mathbf{y}}_{t}$ achieves the lowest loss (Figure˜6).
@@ 4.3 Schedule-Free Methods at the Edge of Stability
We return to the toy model from Section˜4.1 to further analyze the optimization dynamics of the ${\mathbf{y}}_{t}$ iterates. Notably, the ${\mathbf{y}}_{t}$ sequence exhibits a period-two oscillation centered along the river (Figure˜5). This behavior resembles the dynamics of full-batch GD at the Edge of Stability (EoS), where iterates oscillate along sharp directions while remaining globally stable. The EoS phenomenon was first identified empirically by Cohen et al. (2021) and has since been studied theoretically (Arora et al., 2022; Wang et al., 2022; Ahn et al., 2023; Damian et al., 2023; Song and Yun, 2023; Zhu et al., 2023). A key feature of this regime is that the sharpness—measured by the largest Hessian eigenvalue—stabilizes near the threshold $2/\gamma$.
@@
Notation.  Let $f({\mathbf{w}})$ be the objective and $\mathbf{H}({\mathbf{w}})$ its Hessian at $\mathbf{w}$}. For brevity, write $\mathbf{H}_{t}\coloneqq\mathbf{H}({\mathbf{w}}_{t})$. The largest eigenvalue $\lambda_{1}(\mathbf{H})$ is called the sharpness, and for a given preconditioner $\mathbf{P}$, $\lambda_{1}(\mathbf{P}^{-1}\mathbf{H})$ is called the preconditioned sharpness.
We first study the stability of Schedule-Free GD (SF-GD) on quadratic objectives, which serve as local Taylor approximations to neural network training losses. SF-GD is defined by (SF) with $\Delta_{t}\triangleq\nabla f({\mathbf{y}}_{t})$.
@@
Next, we extend the analysis to SF-PrecondGDW, defined by (SF) with $\Delta_{t}\triangleq\mathbf{P}^{-1}\nabla f({\mathbf{y}}_{t})+\lambda{\mathbf{y}}_{t}$, where $\mathbf{P}$ is a fixed, symmetric, and positive definite preconditioner, and $\lambda$ denotes the weight decay coefficient. This setting parallels the analysis of Adaptive Edge of Stability in Cohen et al. (2022).
@@
Proofs of Propositions˜4.1 and 4.2 are deferred to Section˜D.1.
SF-AdamW can be viewed as SF-PrecondGDW with with a slowly varying diagonal preconditioner $\mathbf{P}_{t}$. Hence, its stability is governed by the preconditioned sharpness $\lambda_{1}(\mathbf{P}_{t}^{-1}\mathbf{H}_{t})$. As shown in Figure˜7, the preconditioned sharpness at ${\mathbf{y}}_{t}$ iterates equilibrates near the stability threshold in both the toy model and CIFAR-10 experiments—exhibiting a typical EoS behavior.
Figure 7: The ${\mathbf{y}}_{t}$ iterates of Schedule-Free methods operate at the Edge of Stability. Plots of (preconditioned) sharpness during full-batch training; dashed lines indicate stability thresholds. Left: Toy model trained using SF-AdamW with $(\beta_{1},\beta_{2})=(0.9,0.99)$. Middle, Right: Fully connected network trained on a 5k subset of CIFAR-10 using SF-GD with $\beta=0.9$ and SF-AdamW with $(\beta_{1},\beta_{2})=(0.95,0.99)$; (preconditioned) sharpness is evaluated at the $\mathbf{y}_{t}$ iterates (Figure˜7).